OpenStudy (h0pe):

Find the value of \[x = 1 + \cfrac{1}{2 + \cfrac{1}{2 + \cfrac{1}{2 + \cfrac{1}{2 + \ddots}}}}.\]

OpenStudy (amoodarya):

hint : use this substitution x-1=y so \[y=2+\frac{1}{2+\frac{1}{2+}...}\]

OpenStudy (h0pe):

How does that get me anywhere....?

ganeshie8 (ganeshie8):

after that, you just need to solve \[\large y = 2+\frac{1}{y}\]

OpenStudy (h0pe):

so it would get me to \[y=\sqrt{2}\]

OpenStudy (h0pe):

oh

ganeshie8 (ganeshie8):

because the entire bottom part is same as "y" : \[y=2+\frac{1}{\color{red}{2+\frac{1}{2+}...}} = 2+\frac{1}{\color{red}{y}}\]

OpenStudy (h0pe):

Oh... I understand now. Thank you

ganeshie8 (ganeshie8):

you should get \(y = \sqrt{2}-1\) thus \(x = y+1 = \sqrt{2}\)

OpenStudy (amoodarya):

\[y=2+\frac{1}{y}\\y^2=2y+1\\y^2-2y-1=0\\(y-1)^2-2=0\\(y-1)=\pm \sqrt{2}\\y=1+\sqrt{2}\]because y>0